# DISRetrieval: Harnessing Discourse Structure for Long Document Retrieval

### v1

@@ 3.2 Node Representation
/* Stage 1: Multi-level Evidence Retrieval */
@@
  // Transform query to embedding space
$\text{scores}\leftarrow\{\}$, $E\leftarrow\{\}$ ;
for node $v\in T$ do
       $\text{scores}[v]\leftarrow\text{cosine}(\mathbf{e_{q}},\mathbf{e_{v}})$ ;
        // Compute similarity scores
end for
$V_{\text{ranked}}\leftarrow\text{sort}(\text{scores},\text{descending}=\text{True})$ ;
/* Stage 2: Structure-aware Evidence Selection */
[…]
  // Track selected leaves
@@
       if $v$ is leaf node then
[…]
                   $E\leftarrow E\cup\{v\}$ ;
                    // Direct selection of relevant leaves
[…]
             end if
@@
             $L_{v}\leftarrow\text{get\_leaves}(v)$ ;
              // Get all leaves from subtree
[…]
              // Remove already selected leaves
             if $|L^{\prime}_{v}|>0$ then
                   $L_{\text{topk}}\leftarrow\text{top\_k}(L^{\prime}_{v},\text{scores},k)$ ;
                    // Leaf TopK Selection
                   $E\leftarrow E\cup L_{\text{topk}}$ ;
[…]
@@
Algorithm 1 Hierarchical Evidence Retrieval and Selection
LLM Enhancement.
The document-level discourse tree $T_{\text{doc}}$ contains only discourse relationships at internal nodes. To obtain semantic representations for these nodes, we apply bottom-up LLM enhancement following Equation 2, similar to the paragraph-level process. This enhancement ensures that all nodes in the tree have meaningful semantic representations.
Tree Integration.
After enhancing the document-level tree $T_{\text{doc}}$, we perform tree integration to combine it with all paragraph-level trees $T_{i}$ into a unified discourse structure. The integration process is straightforward: each leaf node in $T_{\text{doc}}$ is replaced with its corresponding paragraph-level discourse tree $T_{i}$, resulting in our final integrated tree $T_{\text{D}}$. This hierarchical structure effectively preserves both local sentence relationships within paragraphs and global discourse connections across the document.
Node Encoding.
To enable efficient retrieval, we transform the discourse tree into an embedding tree structure using an off-the-shelf pre-trained sentence encoder. The resulting embedding tree preserves both the hierarchical structure and semantic information of the original discourse tree, making it suitable for downstream semantic matching tasks.
3.3 Hierarchical Evidence Retrieval and Selection
Our hierarchical evidence retrieval process introduces a novel structure-aware selection mechanism that leverages both local and global discourse information. Our approach combines:
Direct selection of highly relevant leaf nodes Indirect selection through relevant internal nodes Automatic redundancy elimination via tree structure
Specifically, our algorithm first encodes the input query and computes similarity scores with all tree nodes. Then it processes nodes in order of decreasing relevance: leaf nodes are selected directly if highly relevant, while for internal nodes, we identify and select the most relevant unselected leaves from their subtrees. This process continues until we reach the desired evidence set size, ensuring both local relevance and structural coherence. The complete retrieval process is illustrated in Figure 2, demonstrating how evidence is gathered from different levels of the discourse tree. For implementation details, we refer readers to Algorithm 1.
@@ 4.1 Experimental Setup
We evaluate our approach on two challenging long document QA datasets: QASPER [12] for research paper understanding and QuALITY [30] for long-form reading comprehension. For QASPER, we use F1-Match for QA performance and token-level F1/Recall for retrieval evaluation. For QuALITY, we report accuracy as the primary metric. To ensure fair comparison, we fix the total length of retrieved context across all methods.
@@
  • RAPTOR constructs a semantic tree through recursive embedding, clustering, and summarization, with retrieval performed on a collapsed tree structure following Sarthi et al. [34].
@@
For discourse analysis, we construct our sentence-level RST training data based on RST-DT [6]. We then train our discourse parser following Yu et al. [41], adopting XLNet-base-cased as the pre-trained language model backbone. All summarization tasks, including both our approach (as defined in Equation 2) and RAPTOR baseline, utilize Llama3.1-8B-Instruct for consistency. We set the temperature parameter $\tau$ in Equation 2 to 0 for QASPER and 100 for QuALITY in our LLM enhancement process. For sentence encoder (Section §3.2), we experiment with both Sentence-BERT111multi-qa-mpnet-base-cos-v1, https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 and the latest embedding model of OpenAI [^2] OpenAI text-embedding-3-large to evaluate the effectiveness of our method across different representation spaces. The Top-K parameter in leaf selection is set to 5 based on preliminary experiments (detailed in Appendix E.2). For generation models, we employ UnifiedQA-3B, GPT-4.1-mini and Deepseek-v3 to demonstrate the robustness of our retrieval improvements across different generation capabilities. More details are provided in Appendix D.
@@ 4.2 Main Results
We conduct extensive experiments across different context lengths (200-400 words) and embedding models (SBERT and OpenAI). The results are shown in Table 1, comparing our approach with baselines on QASPER and QuALITY datasets.
@@
    DISRetrieval 37.44 40.01 40.28 42.10 41.86 43.21 68.36 72.53 71.62 75.50 73.68 77.42
Our experimental results demonstrate three key findings:
First, DISRetrieval consistently outperforms all baseline methods across different experimental settings. With 400-word contexts utilizing UnifiedQA-3B, it achieves substantial improvements over the flatten-sentence baseline (40.25% vs. 37.99% F1-Match on QASPER; 58.91% vs. 55.27% accuracy on QuALITY). These performance advantages remain robust and consistent across various generation models (UnifiedQA-3B, GPT-4.1-mini, and Deepseek-v3).
Second, compared to the tree-based RAPTOR baseline, DISRetrieval shows clear advantages. Using OpenAI embeddings with 300-word context, our method achieves 40.24% F1-Match versus RAPTOR of 39.96% on QASPER, demonstrating that incorporating linguistic discourse structure is more effective than purely relying on semantic-based clustering for document organization.
Finally, ablation studies with the Bisection variant highlight the importance of discourse structure in our approach. While Bisection improves over flatten-based methods through hierarchical modeling, it consistently underperforms the full DISRetrieval implementation across all settings. This empirically confirms that the integration of linguistic-informed discourse structure provides substantial benefits beyond simple hierarchical organization for long document understanding.
@@
We evaluate the retrieval effectiveness of DISRetrieval on the QASPER dataset using token-level F1 and Recall metrics. The results in Table 2 demonstrate three key findings:
@@
Table 2: Token-level F1 and Recall scores for retrieved contexts on the QASPER dataset across different methods using SBERT and OpenAI embedding model. Best results are bolded and runner-ups are underlined.
Method 200 (SBERT) 200 (OpenAI) 300 (SBERT) 300 (OpenAI) 400 (SBERT) 400 (OpenAI)
@@
   DISRetrieval 28.45 61.28 30.29 64.92 24.67 68.25 26.05 71.62 21.70 73.09 22.63 75.77
@@ 4.3 Discussions
RQ1: Is the hierarchical retrieval strategy effective? Fig 4
In Section 3.3, we propose a carefully designed hierarchical retrieval strategy composed of multiple sub-steps and associated methods. To assess the contribution of each component, we conduct ablation experiments on the QuALITY dataset to validate the effectiveness of each component in our hierarchical retrieval strategy. As shown in Figure 4, we compare five variants: leaf-only baseline, summary-based retrieval, all filtered-leaves, Top-K with ranking order, and our final Top-K with original order.
The results reveal three key insights: (1) First, using summaries of intermediate nodes performs worse than the leaf baseline, indicating that preserving original text details is crucial. (2) Second, while using all filtered leaves shows slight improvements, it is less effective than selective Top-K methods due to noise from irrelevant content. (3) Most importantly, our Top-K origin method, which preserves the natural document flow, consistently achieves the best performance across all settings. This advantage becomes more pronounced with longer contexts and better embeddings, demonstrating that maintaining document structure is essential for effective retrieval.
@@
Table 3: Effect of different LLMs for node text enhancement on QASPER dataset retrieval performance (token-level F1 and Recall) across varying context lengths (200, 300, and 400 words).
@@
RQ2: Does the scale of LLMs affect the quality of the discourse-aware tree structures in node text enhancement? Tab 4.3
We investigate whether model scale affects the quality of discourse-aware tree construction by comparing different LLMs. As shown in Table 4.3, smaller models like Llama-3.1-8B, Qwen2.5-7B, and Mistral-7B achieve comparable performance to larger models, with differences under 0.5%. Notably, Llama-3.1-8B achieves the best recall at 400-word context length. These results indicate that the effectiveness of our discourse-aware retrieval does not heavily depend on LLM scale. This finding is particularly valuable as it enables flexible model selection without requiring extremely large models, making our approach more practical and cost-effective.
@@
RQ3: Does the capability of the discourse parser have a significant impact? Fig 5
@@ 5 Conclusion
In this paper, we presented DISRetrieval, a discourse-aware hierarchical retrieval framework that advances long document understanding through three key innovations: (1) a multi-granularity adaptive discourse framework based on RST, (2) an LLM-enhanced node representation technique, and (3) a structure-aware evidence selection mechanism. Our evaluation on QASPER and QuALITY benchmarks demonstrates significant improvements over existing methods. The consistent performance advantages across varying document lengths and embedding models validate the effectiveness of our discourse-aware approach. Through detailed ablation studies, we identified that the combination of discourse-aware tree construction and hierarchical retrieval mechanism contributes most significantly to the performance gains, highlighting the crucial role of linguistic structure in document understanding. Looking forward, we envision three promising directions: (1) extending our framework to cross-lingual document retrieval, (2) incorporating more sophisticated discourse theories, and (3) exploring dynamic discourse structure adaptation for different domains. Our work establishes a new paradigm for leveraging linguistic structures in information retrieval systems, paving the way for more effective long-form content understanding.
Appendix A Limitation and Future Work
We here discuss the limitations and future work of DISRetrieval.
First, while our discourse-aware tree construction method’s performance is inherently bounded by the quality of RST parsing, our current approach demonstrates significant improvements even with a straightforward adaptation from EDU-level to sentence-level RST structures. This success validates the effectiveness of incorporating discourse information in document retrieval. However, we believe there is still substantial room for enhancement. In future work, we plan to explore more sophisticated discourse analysis techniques and develop specialized tree construction methods that can better capture the hierarchical relationships in long documents.
Second, although we have considered computational efficiency by implementing an adaptive summarization strategy with a length threshold $\tau$ to determine when LLM processing is necessary, this approach is relatively simple. While effective, it may not always achieve an optimal balance between computational cost and representation quality. Future work will investigate more sophisticated methods for managing computational resources, such as dynamic thresholding based on content complexity, hierarchical position-aware summarization, and more efficient LLM architectures specifically designed for tree-structured document processing.
Third, while our evaluation on QASPER and QuALITY datasets demonstrates significant improvements in traditional retrieval metrics, the assessment of long document understanding could benefit from more diverse evaluation criteria. Current metrics may not fully capture the nuanced aspects of discourse-aware retrieval, such as structural coherence preservation and hierarchical information flow. We plan to develop more comprehensive evaluation frameworks that consider multiple dimensions of retrieval quality, including discourse coherence, semantic relevance, and structural preservation at different granularity levels.
These limitations point to promising directions for future research while reinforcing the solid foundation and significant contributions of our current work. We believe addressing these aspects will further advance the field of discourse-aware document retrieval and enhance its practical applications.
Appendix B Potential Impact
The current approaches to long document retrieval often struggle with maintaining semantic coherence and structural integrity, primarily due to the lack of discourse-aware document modeling. Our work, DISRetrieval, addresses this fundamental challenge by introducing a novel framework that leverages discourse structure for enhanced document understanding and retrieval.
The impact of our work extends beyond mere performance improvements in retrieval metrics. By demonstrating the effectiveness of discourse-aware hierarchical modeling, we provide the research community with a new paradigm for approaching long document understanding tasks. Our framework shows that incorporating discourse structure not only enhances retrieval accuracy but also preserves the natural organization of information, which is crucial for human comprehension.
Moreover, our LLM-enhanced node representation technique introduces a novel way of combining traditional discourse analysis with modern LLMs. This hybrid approach could inspire similar innovations in other NLP tasks where maintaining structural relationships while leveraging powerful language models is essential. The adaptive summarization strategy we proposed, although currently implemented with a simple threshold mechanism, opens up new research directions for efficient document processing with LLMs.
Following this line of research, similar discourse-aware approaches could be applied to various challenging tasks in document understanding, such as long-form question answering, document summarization, and information extraction. Our work provides the community with a new perspective on handling long documents by emphasizing the importance of discourse structure, potentially influencing how future systems approach the challenge of processing and understanding lengthy textual content.
Furthermore, the success of DISRetrieval in bridging the gap between document structure and semantic understanding suggests promising applications in real-world scenarios, particularly in domains where maintaining the logical flow of information is crucial, such as legal document analysis, scientific literature review, and educational content processing.
Appendix C Extension of Technical Details

### v2

@@ 3.3 Structure-Guided Evidence Retrieval
/* Compute node similarities and rank */
@@
$\text{scores}[v]\leftarrow\text{cosine}(\mathbf{e_{q}},\mathbf{e_{v}})$ for all $v\in T$ ;
$V_{\text{ranked}}\leftarrow\text{sort}(\text{scores},\text{descending})$ ;
/* Structure-aware selection */
$E\leftarrow\{\}$, $\text{used}\leftarrow\{\}$ ;
@@
 if $v$ is leaf and $v\notin\text{used}$ then
    $E\leftarrow E\cup\{v\}$, $\text{used}\leftarrow\text{used}\cup\{v\}$ ;
@@
    $L\leftarrow\text{leaves}(v)\setminus\text{used}$ ;
    // Unused leaves in subtree
    if $L\neq\emptyset$ then
       $L_{k}\leftarrow\text{top\_k}(L,\text{scores},k)$ ;
       $E\leftarrow E\cup L_{k}$, $\text{used}\leftarrow\text{used}\cup L_{k}$ ;
@@
Algorithm 1 Structure-Guided Evidence Retrieval
@@ 4.1 Experimental Setup
We evaluate our approach on three challenging long document QA datasets: QASPER (Dasigi et al., 2021) for research paper understanding (averaging 4170 words), QuALITY (Pang et al., 2022) for long-form reading comprehension (averaging 5022 words), and NarrativeQA (Kočiský et al., 2018) for QA over exceptionally long narrative documents (averaging 51,372 words). For QASPER, we use F1-Match for QA performance and token-level F1/Recall for retrieval evaluation. For QuALITY, we report accuracy as the primary metric. And for NarrativeQA, we measure performance using BLEU(B-1), ROUGE(R-L) and METEOR(M) metrics. To ensure fair comparison, we fix the total length of retrieved context across all methods.
@@
  • RAPTOR constructs a semantic tree through recursive embedding, clustering, and summarization, with retrieval performed on a collapsed tree structure following Sarthi et al. (2024).
@@
For discourse analysis, we construct our sentence-level RST training data based on RST-DT (Carlson et al., 2001). We then train our discourse parser following Yu et al. (2022b), adopting XLNet-base-cased as the pre-trained language model backbone. All summarization tasks, including both our approach (as defined in Equation 2) and RAPTOR baseline, utilize Llama3.1-8B-Instruct for consistency. We set the threshold parameter $\tau$ in Equation 2 to 0 for QASPER and 50 for QuALITY and NarrativeQA (detailed in Appendix C.4). For sentence encoder (Section §3.2), we experiment with both Sentence-BERT111multi-qa-mpnet-base-cos-v1, https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 and the latest embedding model of OpenAI [^2] OpenAI text-embedding-3-large to evaluate the effectiveness of our method across different representation spaces. The Top-K parameter in leaf selection is set to 5 based on preliminary experiments (detailed in Appendix C.3). For generation models, we employ UnifiedQA-3B, GPT-4.1-mini and Deepseek-v3 to demonstrate the robustness of our retrieval improvements across different generation capabilities. More details are provided in Appendix B.
@@ 4.2 Main Results
We evaluate our approach across different context lengths (200-400 words) and embedding models (SBERT and OpenAI) on QASPER and QuALITY datasets. Table 1 presents the comparative results. The results reveal three key insights:
@@
    DISRetrieval 38.34 40.03 40.33 42.10 41.61 43.30 68.94 72.96 72.39 76.70 73.78 78.00
Consistent superiority across settings. DISRetrieval outperforms all baselines regardless of context length, embedding model, or generation architecture (UnifiedQA-3B, GPT-4.1-mini, and Deepseek-v3). For instance, with 400-word contexts and UnifiedQA-3B, we achieve +2.58% F1-Match on QASPER and +4.37% accuracy on QuALITY over the flatten-sentence baseline.
Discourse structure surpasses semantic clustering. Compared to RAPTOR’s semantic clustering approach, our discourse-aware method demonstrates clear advantages (40.24% vs. 39.96% F1-Match on QASPER with OpenAI embeddings), confirming that linguistic discourse structure provides more principled document organization than purely semantic-based methods.
Linguistic discourse essential for hierarchical modeling. The Bisection ablation validates our core hypothesis: while hierarchical organization helps (Bisection > flatten baselines), incorporating discourse structure is crucial (DISRetrieval > Bisection consistently), providing substantial benefits beyond simple tree-based document modeling.
@@
We evaluate DISRetrieval’s retrieval effectiveness on QASPER using token-level F1 and Recall metrics. The Retrieval Result in Table 2 reveals three key findings:
@@
Retrieval Result 200 (SBERT) 200 (OpenAI) 300 (SBERT) 300 (OpenAI) 400 (SBERT) 400 (OpenAI)
@@
  DISRetrieval 28.66 61.76 30.57 65.51 24.74 68.30 26.20 72.20 21.68 72.92 22.76 76.03
Generation Result Full Document + Retrieved Evidence Full Document Gold Evidence
200 (SBERT) 200 (OpenAI) 300 (SBERT) 300 (OpenAI) 400 (SBERT) 400 (OpenAI) Avg. 4170 wrods Avg. 129 words
F1 / % 49.56 (+0.75) 49.83 (+1.02) 49.63 (+0.82) 49.72 (+0.91) 49.71 (+0.90) 49.56 (+0.75) 48.81 50.71 (+1.90)
Table 2: Retrieval and generation results (token-level F1 and Recall) on the QASPER dataset using SBERT and OpenAI Embedding as embedding models and Llama3.1-8B-Instruct as generation model. Best results are bolded and runner-ups are underlined.
BLEU / % ROUGE / % METEOR / %
flatten-chunk 24.24 28.42 19.70
flatten-sentence 21.53 28.22 18.33
RAPTOR 25.05 30.24 20.92
Ours
Bisection 24.71 28.85 20.41
DISRetrieval 25.59 30.26 21.21
Table 3: Performance on the NarrativeQA dataset. Best results are bolded and runner-ups are underlined.
@@ 4.3 Discussions
RQ1: Is DISRetrieval effective for handling extremely long documents? Tab 3
We evaluate our method on NarrativeQA, which contains exceptionally long documents (average: 51,372 words, maximum: 346,902 words) that exceed most generative models’ context limits. DISRetrieval achieves superior performance across all metrics, outperforming the widely-used flatten-chunk baseline by +1.35% BLEU, +1.84% ROUGE, and +1.51% METEOR. Notably, both DISRetrieval and RAPTOR substantially outperform flatten-based methods, confirming that structured approaches are more effective than flat for extremely long document retrieval. However, DISRetrieval’s discourse-aware structure provides consistent advantages over RAPTOR’s semantic clustering approach.
RQ2: Is precise evidence retrieval essential for effective question answering? Tab 2
To evaluate the importance of precise evidence retrieval, we compare three context settings using Llama3.1-8B-Instruct: (i) golden evidence, (ii) full document, and (iii) full document augmented with retrieved evidence (highlighted with [EVIDENCE] … [/EVIDENCE] markers). Golden evidence achieves the highest F1 score of 50.71% with only 129 words on average, significantly outperforming the full document approach (48.81% with 4,170 words). This demonstrates that concise yet accurate retrieval is more effective than using substantially larger context. Furthermore, augmenting the full document with our retrieved evidence yields consistent improvements of 0.75-1.02% across all settings, confirming that precise retrieval enhances performance even when combined with complete document access. These results underscore that effective long-document QA systems should prioritize retrieval quality over context quantity.
@@
RQ3: Is the hierarchical retrieval strategy effective? Fig 3
We conduct ablation experiments on the QuALITY dataset to evaluate our hierarchical retrieval strategy. We compare five variants: leaf-only baseline, summary-based retrieval, all filtered-leaves, Top-K with ranking order, and our Top-K with original order. Three key findings emerge: (1) Summary-based retrieval underperforms the leaf baseline, confirming that preserving original text details is crucial for effective retrieval. (2) While all filtered-leaves shows marginal improvements, selective Top-K methods are superior due to reduced noise from irrelevant content. (3) Our Top-K origin method consistently achieves the best performance by preserving natural document flow. This advantage intensifies with longer contexts and stronger embeddings, validating that maintaining document structure is essential for hierarchical retrieval.
RQ4: Does the scale of LLMs affect the quality of the discourse-aware tree structures in node text enhancement? Tab 4
We investigate whether model scale affects the quality of discourse-aware tree construction by comparing different LLMs. As shown in Table 4, smaller models like Llama-3.1-8B, Qwen2.5-7B, and Mistral-7B achieve comparable performance to larger models, with differences under 0.5%. Notably, Llama-3.1-8B achieves the best recall at 400-word context length. These results indicate that the effectiveness of our discourse-aware retrieval does not heavily depend on LLM scale. This finding is particularly valuable as it enables flexible model selection without requiring extremely large models, making our approach more practical and cost-effective.
@@
Table 4: Effect of different LLMs for node text enhancement on QASPER dataset retrieval performance (token-level F1 and Recall) across varying context lengths.
RQ5: Does the capability of the discourse parser have a significant impact? Fig 4
@@ 5 Conclusion
In this paper, we presented DISRetrieval, a discourse-aware hierarchical retrieval framework that systematically incorporates rhetorical structure theory into long document question answering. Through three key innovations—sentence-level discourse parsing adapted for lengthy texts, LLM-enhanced node representations, and structure-aware evidence selection—our approach demonstrates substantial improvements over existing methods across multiple challenging benchmarks. Comprehensive experiments on QASPER, QuALITY, and NarrativeQA demonstrate consistent performance improvements across varying context lengths, embedding models, and generation architectures. Our ablation studies confirm that discourse structure provides more principled document organization than semantic clustering approaches, with the framework naturally adapting to different document types through multi-granularity retrieval. Beyond performance improvements, this work demonstrates that linguistically-grounded approaches remain valuable in the era of LLMs, offering principled alternatives to purely data-driven methods. Our framework opens new directions for incorporating structured linguistic knowledge into neural retrieval systems, with potential applications in legal document analysis, scientific literature review, and educational content processing.
Limitations
While DISRetrieval demonstrates substantial improvements across multiple datasets and architectures, we acknowledge several limitations that point to promising future directions. First, our discourse parser’s performance inherently bounds the overall system effectiveness, as shown in Figure 4 where parser quality directly impacts downstream performance. However, our results demonstrate that even with a straightforward EDU-to-sentence adaptation, significant gains are achieved, validating the core hypothesis that discourse structure benefits long document retrieval. Future work will explore more sophisticated discourse parsing techniques, including domain-adaptive parsers and multi-granularity discourse analysis. Second, our adaptive summarization strategy using threshold $\tau$ is relatively simple but proves effective across diverse document types ($\tau$=0 for academic papers, $\tau$=50 for narratives). While this approach successfully balances computational efficiency with representation quality, more sophisticated dynamic thresholding based on content complexity and hierarchical position could further optimize the trade-off. Third, current evaluation metrics, though comprehensive across three challenging datasets, may not fully capture the nuanced benefits of discourse-aware retrieval, such as structural coherence preservation and hierarchical information flow. We plan to develop evaluation frameworks that better assess discourse-aware retrieval quality across multiple dimensions.
Ethical Statement
This research adheres to ethical principles in natural language processing research and does not raise significant ethical concerns. Our work focuses on improving document retrieval and question answering systems through discourse-aware hierarchical modeling, which has potential positive societal impacts by enhancing information access and comprehension. The datasets used in our experiments (QASPER, QuALITY, and NarrativeQA) are publicly available benchmark datasets that have been previously vetted by the research community and do not contain sensitive personal information or harmful content. Our discourse parsing approach operates on linguistic structures rather than content semantics, minimizing risks of bias amplification or misuse. The proposed DISRetrieval framework is designed as a general-purpose retrieval enhancement technique that can benefit various applications requiring long document understanding, such as academic research assistance, educational content processing, and legal document analysis. We acknowledge that like any information retrieval system, our method could potentially be misused if applied to spread misinformation or manipulate access to information, but such concerns are not specific to our approach and apply broadly to information retrieval technologies. We encourage responsible deployment of our system with appropriate safeguards and human oversight, particularly in high-stakes applications. All experiments were conducted using publicly available computational resources and open-source tools, ensuring reproducibility and transparency in our research process.
Appendix A Extension of Technical Details